Activity-Based Costing: Selling and Administrative Expenses

Jungle Junior Company manufactures and sells outdoor play equipment. Jungle Junior uses activity-based costing to determine the cost of the sales order processing and the customer return activity. The sales order processing activity has an activity rate of $20 per sales order, and the customer return activity has an activity rate of $100 per return. Jungle Junior sold 2,500 swing sets, which consisted of 750 orders and 80 returns.

a. Determine the total sales order processing and customer return activity cost for swing sets.
$

b. Determine the per-unit sales order processing and customer return activity cost for swing sets. Round your answer to the nearest cent.
$ per unit
Answer #1

a) Total Sales order processing cost = 750*20 = 15000

Customer return activity cost = 80*100 = 8000

b) Per unit = (15000+8000)/2500 = 9.20